OpenStudy (anonymous):

Factor. (x+5)^3-(2x+1)^3

hartnn (hartnn):

\(\Large a^3-b^3= (a-b)(a^2+ab+b^2)\)

hartnn (hartnn):

here, a = x+5, b= 2x+1

hartnn (hartnn):

just plug this in the formula! what u get ?

OpenStudy (anonymous):

Not sure, could you just show me? It'll be easier for me to understand, then I'll try the same with a different question

hartnn (hartnn):